# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ 5 Conclusion
We note that our approach is quite general and is agnostic to the exact nature of the LLM-based approach used. Thus, any of the LLM-based approaches discussed in Section 3 can be used in the first step of our approach. This is important since the LLM models and fine-tuning approaches continue to advance quickly; so the plug-and-play nature of our approach extends well to this setting. In addition, when considering personalization and accessing contextual information (e.g., readers’ features), we can incorporate this extra information as inputs to the LLM-based CTR prediction model and integrate it into contextual bandit algorithms. Specifically, we provide two variants of LOLA for this purpose in Web Appendix B.
@@
In conclusion, our study introduces LOLA, a novel framework that combines the predictive power of LLMs with the adaptive efficiency of online learning algorithms to enhance content experimentation. By leveraging LLM-based prediction models as priors, our approach minimizes regret in real time, leading to significant improvements in total reward/clicks. Furthermore, through the numerical analysis based on a large-scale dataset from Upworthy, we demonstrate that LOLA outperforms traditional A/B testing and pure online learning algorithms, particularly in settings with limited experimental traffic or a large number of arms. These conditions are prevalent in the current content industry, where platform users can easily create and share their content, such as on platforms like X and TikTok. Our numerical results also indicate that while LLMs provide valuable predictions, one cannot fully trust the prediction results solely from LLMs since experimentation-based approaches still remain the gold standard and can be leveraged to correct the errors in LLMs.
From a managerial perspective, LOLA offers a versatile solution applicable to various scenarios where firms need to decide on content display, such as news articles, headlines, advertisements, or website content. Utilizing open-source LLM models, this method is cost-effective and can be easily deployed across different tasks once LLM models are fine-tuned on relevant datasets. The fine-tuning process is cost-efficient and can be performed on entry-level GPUs, making advanced LLM-based techniques accessible to firms with limited budgets. Additionally, we document that prompt-based approaches, which have been the standard method in recent marketing literature, perform poorly in this content attractiveness comparison task, underscoring the importance of the fine-tuning approach.
Finally, we highlight that LOLA can be readily adapted to other measures beyond clicks. The click-based metric, while widely used by Upworthy a decade ago, has raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity. The industry is rapidly evolving, with Upworthy now focusing more on content quality and humanity, although clicks still remain an important measure; see more detailed discussion in (Upworthy2023). Our LOLA framework can easily accommodate these changes by adjusting the reward notion in the algorithm to align with new business goals, demonstrating its flexibility and relevance in a dynamic industry.
@@
Appendix A Supplementary Materials for Section 3
A.1 Pairwise Comparison Between GPT Prompts
A.2 Implementation Details of MLP Classification Models
Hyperparameters of the network structure. We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer. This configuration is widely used in practice. To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$. Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer. This ratio here is in $\{2,4,8\}$. Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer. This is because the final output layer has only one dimension for ranking the most attractive headline. We incorporate the dropout after each hidden layer to mitigate the overfitting. Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer. In this work, we set the dropout rate in $\{0.3,0.5\}$. All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
Hyperparameters of the training procedure. The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing. To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch. Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence. We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$. Additionally, to mitigate overfitting, we apply $L_{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$. We use binary cross-entropy as the loss function and employ stochastic gradient descent for training. We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
A.3 Results of DistilBERT LoRA Fine-Tuning
One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model. An illustration of the structure of the DistilBERT model combined with a classification head is shown in Figure A3. By default, DistilBERT can handle two headlines together as input. The format of the input is: [CLS] Headline 1 [EOS] Headline 2 [EOS]. The [CLS] token is a special token used to get the output of the model for classification, and the output of the last Transformer layer corresponding to this token is used as the input to the classification head. The [EOS] token is used to separate the two sentences.
Another difference is that DistilBERT is a smaller model compared to Llama-3-8b. DistilBERT has around 66 million parameters, which is only 0.94% of the parameters of Llama-3-8b. Due to the smaller model size, we are able to conduct full parameter updates during the fine-tuning process without using LoRA. We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in Section 3.3.3. The accuracy of DistilBERT LoRA fine-tuning on the same test set is around 75%. More details can be found in the code repository.
Appendix B Algorithms with Contextual Information
For the completeness of the paper, we introduce two additional LOLA algorithms in this appendix, which leverage contextual information.
For linear contextual bandits, we assume that at each round, we first observe the context vector $x_{t,k}\in\mathbb{R}^{d}$ for each arm $k\in[K]$. This context information can include LLM embedding vectors and user information, among other features. We denote by $r_{t,k}$ the random reward of arm $k$ at round $t$. After pulling one arm at each round, only the reward of that arm, $r_{t}$, is revealed. We make the linear realizability assumption that the expected reward is linear in the context vectors, which means $\mathbb{E}[r_{t,k}|x_{t,k}]=x_{t,k}^{\top}\theta^{*}$. The weight vector $\theta^{*}\in\mathbb{R}^{d}$ is unknown and needs to be learned. Additionally, we assume access to a historical dataset of size $M$, denoted by $\{(x_{t,k},r_{t})\}_{t=-M}^{-1}$. With these definitions, we are ready to present our proposed Algorithm 2, named LLM-Assisted Linear Contextual Bandit. The LLM-LinUCB algorithm extends LinUCB (chu2011contextual). Under the linear realizability assumption, this algorithm achieves asymptotically optimal regret, matching the lower bound up to logarithmic factors.
Algorithm 2 LLM-Assisted Linear Contextual Bandit (LLM-LinUCB)
1:LLM Training Phase: Calculate $A\leftarrow I_{d}+\sum_{t=-M}^{-1}x_{t,a_{t}}x_{t,a_{t}}^{\top}$ (The $d$-by-$d$ identity matrix), and $b\leftarrow\sum_{t=-M}^{-1}x_{t,a_{t}}r_{t}$
4:for $t=1$ to $T$ do
5:     Get weight estimator: $\theta_{t}\leftarrow A^{-1}b$
6:     Observe $K$ context vectors, $x_{t,1},x_{t,2},\dots,x_{t,K}\in\mathbb{R}^{d}$, generated through an LLM-based embedding model concatenated with user context information.
7:     for $k=1,2,\dots,K$ do
9:     end for
10:     Choose action $a_{t}=\arg\max_{k\in[K]}\hat{r}_{t,k}$.
11:     Observe reward $r_{t}$.
12:     $A\leftarrow A+x_{t,a_{t}}x_{t,a_{t}}^{\top}$.
13:     $b\leftarrow b+x_{t,a_{t}}r_{t}$.
14:end for
The algorithm proceeds as follows: During the LLM training phase, we prepare the matrix $A$ and vector $b$ to estimate the linear weight vector. We may need an LLM-based embedding model to extract embedding vectors from the original text information. Similar to LLM-2UCBs, we fine-tune the upper bound control parameter $\alpha$ using another fold of data and a bisection search to minimize regret on this fine-tuning dataset. The online learning phase remains the same as the standard LinUCB.
The advantages of LLM-LinUCB are its transparency due to the linear assumption, ease of execution with only one hyperparameter, and theoretical soundness with optimal regret. However, it is restricted to linear forms. There are two issues with this assumption. First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over embedding-based classification. Second, once concatenated with user contexts, it is unclear whether the linear form is a good approximation, necessitating further numerical study.
To alleviate the limitations of the linear assumption, we introduce another LLM-assisted contextual bandit algorithm (Algorithm 3), which extends the FAst Least-squares-regression-oracle CONtextual bandits (FALCON) algorithm proposed by (simchi2022bypassing). This type of contextual bandit algorithm generalizes the linear realizability assumption to a more flexible framework, assuming the true prediction model $f^{*}\in\mathcal{F}$. The functional class $\mathcal{F}$ can be very general and is not limited to linear models. This assumption is reasonable and promising because LLMs have demonstrated their capability across a wide range of tasks. Essentially, $\mathcal{F}$ can be viewed as a class of LLM models, and the goal is to find the true LLM model $f^{*}$ that performs best for our task, i.e., $f^{*}(x_{t,k},k)=\mathbb{E}[r_{t,k}|x_{t,k}]$ for any $k\in[K]$ and $x_{t,k}\in\mathcal{X}$. Under this setting, the context information is not restricted to real numbers and can include more general data, such as text information. We now present the LLM-FALCON algorithm.
4:for epoch $m=1,2,\dots$ do
7:     for round $t=\tau_{m-1}+1,\dots,\tau_{m}$ do
8:         Observe contexts $x_{t,k}\in\mathcal{X}$,$\forall k\in[K]$.
10:         Define
11:         Sample $a_{t}\sim p_{t}(\cdot)$ and observe reward $r_{t}$.
12:     end for
13:end for
After getting an LLM trained on new data, then the key step in this algorithm is using the “inverse proportional to the gap” rule to balance exploration and exploitation. At each round, we first identify the empirically best arm $\hat{a}_{t}$ and then compute the gap in expected reward between this arm and the others. The sampling probability $p_{t}(k)$ is inversely proportional to this gap, meaning that a larger estimated reward $\hat{f}_{m}(x_{t,k},k)$ results in a smaller gap and a higher probability of being pulled. For a more detailed discussion of this algorithm, please refer to (simchi2022bypassing). Our purpose here is to adapt this algorithm in the LLM setting and discuss its pros and cons for completeness.
Appendix C LLM-Assisted Thompson Sampling
For completeness, we also present another version of LOLA that combines LLM and Thompson Sampling. We will update its numerical results in the next version of the manuscript.
Algorithm 4 LLM-Assisted Thompson Sampling (LLM-TS)
4:for $t=1$ to $T$ do
6:     Play the arm $a_{t}=\arg\max_{k\in[K]}\theta_{k}^{t}$.
7:     Observe the payoff $r_{a_{t}}$.
9:end for
Appendix D Survey Results
To illustrate the difficulty of the headline selection task and compare it with LLM-based methods, we conducted a survey among students, asking respondents which headline was catchier.
Specifically, we conducted the survey study among undergraduate students, acknowledging that these participants were less experienced in editing than Upworthy editors and writers. Despite this limitation, we aim to highlight the challenges of this task for humans. It is important to note that all candidate headlines were chosen by Upworthy editors, who were uncertain which one was catchier, prompting the need for further A/B testing.
@@ Appendix D Survey Results
We designed two similar surveys using Qualtrics. The first survey randomly sampled 3,000 pairs from 39,158 headline pairs, differentiated by CTRs at a 0.05 significance level, along with 500 pairs showing no significant differences in CTRs. The second survey utilized the same 3,000 significant samples as the first. In both surveys, students were asked to select the catchier headline from each pair. The introductory text for the surveys was:
@@
After providing identification information, including the last four digits of their phone and student ID, respondents proceeded to a new page featuring 20 headline comparison questions with the prompt, “Please select the headline that you find more attractive,” with only one permissible response. Notably, the first survey included an additional response option: “Both headlines seem equally good.” At the survey’s conclusion, we asked, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
Detailed information and survey results are recorded in Table A1. Because some students did not finish the survey or skipped some comparisons, one can see there are 3,809 responses out of $202\times 20=4,040$. To calculate the accuracy metric, we only focused on the significant pairs, and in the first survey, we ruled out the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys. “Accuracy” measure here is consistent with the LLM tests. We viewed the headline with a higher CTR as the correct answer, and calculated the proportion of correct answers as accuracy. By doing so, one should expect 50% accuracy by random guess since we randomly swapped the order of the headlines to make sure the displayed order does not correlate with the correct answer.
Detailed information and survey results are documented in Table A1. Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$). To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys. Accuracy here aligns with the LLM tests. We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy. Accordingly, one might expect a 50% accuracy rate from random guessing, as we randomized the order of the headlines to ensure that the display order did not correlate with the correct answer.
The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing, with no accuracies significantly deviating from 50% at the 0.05 significance level. Furthermore, accuracies did not vary significantly across different groups based on news reading frequency. We also incorporated two attention checks in the second survey. These findings suggest that the human ability to discern catchier headlines is limited, highlighting the intrinsic difficulty of the task.

### v2

@@ 5 Conclusion
In conclusion, in this paper, we examine if and how firms can leverage the LLMs to enhance content experimentation in digital platforms. First, we examine how well LLMs can predict which content would be more appealing to users. We find that while LLMs provide informative predictions, even the best performing fine-tuned LLMs are unable to perfectly predict which content will be more appealing to users, and match the accuracy/regret of experimentation-based approaches. As such, firms cannot fully trust the prediction results solely from LLMs and replace experimentation.
@@
We introduce LOLA, a novel experimentation framework that combines the predictive power of LLMs with the adaptive efficiency of online learning algorithms to enhance content experimentation. By leveraging LLM-based prediction models as priors, our approach minimizes regret in real time, leading to significant improvements in total reward. Comprehensive numerical experiments based on a large-scale A/B testing dataset demonstrate that LOLA outperforms the traditional A/B tests and pure online learning algorithms.
From a managerial perspective, LOLA offers a versatile solution to a broad set of scenarios where the firm needs to decide which content to display to users. LOLA is particularly valuable under limited traffic conditions that are common in many digital media platforms. This includes cases where content becomes stale quickly (e.g., news articles) or social media platforms like X and TikTok, where users generate a lot of new content regularly, and as a result, the amount of content is high relative to the amount of traffic. While our focal application is in the context of headline selection in the news and media industry, the framework can easily be applied to other problems, such as digital advertising, email marketing, and promotions. Indeed, a large literature in marketing focuses on how to use experiments in conjunction with machine learning methods to identify optimal treatments and/or personalize treatments; see rafieian_yoganarasimhan_2023 for a detailed review. The LOLA approach can be easily used in many of those settings and help with maximizing outcomes of interest while simultaneously reducing the cost of experimentation. Finally, from an implementation perspective, LOLA offers many advantages. Since it can utilize open-source LLM models, it is cost-effective and can be easily deployed across different tasks once LLM models are fine-tuned on relevant datasets. The fine-tuning process itself is cost-efficient and can be performed on entry-level GPUs, making advanced LLM-based techniques accessible to firms with limited budgets. Relying on open-source LLMs also ensures that the firm’s data stays within the firm and is not accessible to the owners of proprietary LLM models (such as Google and OpenAI). Further, our approach is compatible with many bandit-based experimental systems used by large digital firms such as Amazon (fiez_etal_2024), and as such, is easy to integrate and deploy.
@@
Appendix A User Survey
We conduct a survey of undergraduate students from a large public university. The goal of the survey was to examine whether lay human users can accurately identify which headlines would lead to more clicks. Recall that the headlines themselves were created by Upworthy editors, but those editors need an experiment to identify the most appealing headline.
@@ Appendix A User Survey
We design two similar surveys using Qualtrics. The first survey randomly samples 3,000 pairs of headlines from 39,158 headline-pairs that are significantly different from each other (at a 0.05 significance level), along with 500 pairs that have no significant differences in CTRs. The second survey utilizes the same 3,000 significant samples as the first but does not include the insignificant samples.
In both surveys, students were asked to select the catchier headline from a pair of headlines. The introductory text for the surveys was:
@@
Respondents then proceed to a new page featuring 20 headline comparison questions with the prompt, “Please select the headline that you find more attractive,” with only one permissible response. In all the comparisons, we randomly swap the order of the headlines to ensure that the display order does not correlate with the correct answer. We also incorporated two attention checks in the second survey.
The first survey included an additional response option: “Both headlines seem equally good.” At the survey’s conclusion, we also collected data on users’ news consumption habits. Specifically, we ask, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
Detailed information on the surveys and the results from both surveys are shown in Table A1. Note that the total responses received is less than the number of respondents ×$\times 20$ because some students did not finish the survey or skipped some comparisons. For example, for the first survey, the total number of responses 3,809 is less than $202\times 20=4,040$. To calculate the accuracy metric, we only focus on the significant pairs, and in the first survey, we discard the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys. “Accuracy” measure here is consistent with the LLM tests. We treat the headline with a higher CTR as the correct answer and calculate the proportion of correct answers as accuracy. By doing so, one should expect 50% accuracy by random guess since
Detailed information and survey results are documented in Table A1. Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$). To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys. Accuracy here aligns with the LLM tests. We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy. As such, we should expect a 50% accuracy rate from random guessing.
The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing – the accuracy rate in both surveys is not significantly different from 50% at the 0.05 significance level. Furthermore, accuracies did not vary significantly across different groups based on news reading frequency. These findings suggest that the human ability to discern catchier headlines is limited, highlighting the intrinsic difficulty of the task.